# Performance Rating Equilibrium

### v2

@@ 2 Setup
I show that the function TPR has a fixed point $x^{*}\in\mathbb{R}^{n}_{+}$ such that $\text{TPR}(x^{*})=x^{*}$. Define the set $C=\{x\in\mathbb{R}_{+}^{n}\mid x_{i}\leq c,\,i\in N\}$. $C$ is compact, as it is closed and bounded.
First, I show that by the Intermediate Value Theorem, for each $x\in C$ and every player $i$ with $0<m_{i}<k_{i}$, there exists a unique $y_{i}$ that satisfies $m_{i}=g_{i}(y_{i})$, where
$g_{i}(y_{i})=\sum_{j=1}^{k_{i}}\frac{1}{1+10^{\frac{x_{i_{j}}-y_{i}}{400}}}.$
Consider the derivative of each term in the summation with respect to $y_{i}$.
$\frac{dg_{i}}{dy_{i}}=\frac{\ln(10)}{400}\frac{10^{\frac{x_{i_{j}}-y_{i}}{400}}}{\left(1+10^{\frac{x_{i_{j}}-y_{i}}{400}}\right)^{2}}.$ (2)
Since this derivative is strictly positive for each term in the summation, the function $g_{i}(y_{i})$ is strictly increasing in $y_{i}$. As $y_{i}$ approaches $\infty$- ∞, $g_{i}(y_{i})$ tends to 0 because $10^{\frac{x_{i_{j}}-y_{i}}{400}}$ tends to $\infty$. Conversely, as $y_{i}$ approaches +$\infty$+ ∞, $g_{i}(y_{i})$ tends to $k_{i}$ since each term $1/(1+10^{\frac{x_{i_{j}}-y_{i}}{400}})$ approaches 1. Since $g_{i}(y_{i})$ is continuous, strictly increasing, and ranges from 00 to $k_{i}$, by the Intermediate Value Theorem, for any $m_{i}$, there exists a unique $y_{i}$ such that $g_{i}(y_{i})=m_{i}$.
Second, I show that TPR is continuous by applying the Implicit Function Theorem. Define the function $F(y_{i},x)$ as follows.
$F(y_{i},x)=\sum_{j=1}^{k_{i}}\frac{1}{1+10^{\frac{x_{i_{j}}-y_{i}}{400}}}-m_{i}.$
The partial derivative of $F$ with respect to $y_{i}$ is shown in Equation 2, and it is the negative of the partial derivative of $F$ with respect to $x_{i_{j}}$. Since the partial derivatives involve exponential functions, they are continuous with respect to both $y_{i}$ and $x_{i_{j}}$.
Now, calculate $\frac{\partial F}{\partial y_{i}}$:
$\frac{\partial F}{\partial y_{i}}=\sum_{j=1}^{k_{i}}\frac{\ln(10)10^{\frac{x_{i_{j}}-y_{i}}{400}}}{400\left(1+10^{\frac{x_{i_{j}}-y_{i}}{400}}\right)^{2}}.$
Since each term in the summation is strictly positive, $\frac{\partial F}{\partial y_{i}}$ is strictly positive.
By the Implicit Function Theorem, there exists an open neighborhood $U$ containing $x$, and a unique continuously differentiable function $y_{i}=\phi_{i}(x)$ defined on $U$, such that $F(\phi_{i}(x),x)=0$ for all $x\in U$.
Third, I extend the function $\phi_{i}(x)$ to $C$. The collection of sets $\{U_{x}\}_{x\in C}$ forms an open cover of $C$. Since $C$ is compact, there exists a finite subcover $\{U_{1},U_{2},\dots,U_{t}\}$ such that
$C\subseteq\bigcup_{j=1}^{t}U_{j}.$
For each $x\in C$, choose any $U_{j}$ containing $x$ and set $\phi_{i}(x)=\phi_{i}^{(j)}(x)$. For any $x\in U_{j}\cap U_{j^{\prime}}$, uniqueness of the solution $y_{i}$ implies that
$\phi_{i}^{(j)}(x)=\phi_{i}^{(j^{\prime})}(x)=y_{i}.$
Thus, the function $\phi_{i}$ is uniquely defined on the domain $C$. Since each $\phi_{i}^{(j)}$ is continuous on $U_{j}$, $\phi_{i}(x)$ is continuous on $C$. As a result, for every player $i$ with $0<m_{i}<k_{i}$, $\text{TPR}_{i}(x)=\phi_{i}(x)$ is continuous.
Since for every player $i$ with $m_{i}=0$ or $m_{i}=k_{i}$, $\text{CPR}_{i}$ is also a continuous function (Ismail, 2023), the function TPR is a continuous function.
Fourth, by definition, for every $x\in C$ and every $i\in N$, $\text{TPR}_{i}(x)\leq c$. Thus, the function TPR maps $C$ into itself. Moreover, $C$ is convex: let $x,x^{\prime}\in C$ and take any $\lambda\in[0,1]$. For each $i$, $0\leq x_{i}\leq c$ and $0\leq x^{\prime}_{i}\leq c$ implies that $0\leq\lambda x_{i}+(1-\lambda)x^{\prime}_{i}\leq c$. Thus, $(\lambda x+(1-\lambda)x^{\prime})\in C$.
Finally, since the function TPR, when restricted to $C$, is continuous and maps $C$ into itself, by Brouwer’s fixed point theorem, there exists $x^{*}\in C$ such that $\text{TPR}(x^{*})=x^{*}$. ∎
This result extends to other continuous rating functions defined on a convex and compact domain, including the Complete Performance Rating (CPR) and simplified versions of TPR that use average ratings (see, e.g., Ismail, 2023).

### v3

@@ 2 Setup
Fix a tournament $T$ and realized score vector $m=(m_{i})_{i\in N}$. For each player $i\in N$, define the component function $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ by the implicit equation
$\sum_{j\neq i}w_{ij}\,\psi\Big(\frac{f_{i}(x)-x_{j}}{\beta}\Big)=m_{i}.$ (1)
Because $\psi$ is continuous, strictly increasing, and satisfies $\lim_{t\to-\infty}\psi(t)=0$ and $\lim_{t\to\infty}\psi(t)=1$ by Definition 1, the left-hand side of (1) is continuous and strictly increasing in the unknown $z=f_{i}(x)$, with limits 0 and $\sum_{j\neq i}w_{ij}=k_{i}$ as $z\to$-$\infty$ and $z\to\infty$, respectively. Since $m_{i}\in(0,k_{i})$, the intermediate value theorem ensures the existence of a unique solution $z=f_{i}(x)$.
Collect these component functions into the vector function
$f(x):=\big(f_{1}(x),\dots,f_{n}(x)\big)^{\top}:\mathbb{R}^{n}\to\mathbb{R}^{n},$
which we call the canonical $\psi$–performance function.
Now we show the equivalence.
($\Rightarrow$) Suppose $f(x)=x$. Then for each $i\in N$, substituting $f_{i}(x)=x_{i}$ into (1) yields
$\sum_{j\neq i}w_{ij}\,\psi\Big(\frac{x_{i}-x_{j}}{\beta}\Big)=m_{i},$
that is, $E_{i}(x)=m_{i}$. Hence $E(x)=m$.
($\Leftarrow$) Conversely, suppose $E(x)=m$. Then for each $i\in N$,
$\sum_{j\neq i}w_{ij}\,\psi\Big(\frac{x_{i}-x_{j}}{\beta}\Big)=m_{i},$
so $x_{i}$ satisfies (1). Because the solution to (1) is unique, we have $f_{i}(x)=x_{i}$ for all $i$. Therefore $f(x)=x$.
Thus $f(x)=x$ if and only if $E(x)=m$. In particular, a fixed point $x^{*}$ of $f$ satisfies $E(x^{*})=m$, which is exactly the defining condition of a PRE. ∎
In Section 3, we characterize the PRE as the maximizer of a concave potential.
2.1 The standard Elo model
Here, we note that the standard Elo system corresponds to a case in which $\psi$ is a logistic function.
Proposition 1.
The standard Elo win probability function, defined with a scaling factor of 400 and $\beta=1$ as
$\psi_{\text{Elo}}(t)=\frac{1}{1+10^{-t/400}},$ (2)
satisfies the properties of a win probability function as per Definition 1.
Proof.
We verify the required properties for $\psi(t)=\psi_{\text{Elo}}(t)$. Let $C=400$. In terms of domain and range, $\psi(t)$ is defined for all $t\in\mathbb{R}$. Since $10^{-t/C}>0$, we have $1+10^{-t/C}>1$, so $0<\psi(t)<1$.
As for smoothness, $\psi(t)$ is composed of elementary functions and is therefore $C^{1}$.
For monotonicity, we calculate the derivative.
$\psi^{\prime}(t)$ $=\frac{d}{dt}(1+10^{-t/C})^{-1}$
$=-(1+10^{-t/C})^{-2}\cdot(10^{-t/C}\cdot\ln(10)\cdot(-1/C))$
$=\frac{\ln(10)}{C}\frac{10^{-t/C}}{(1+10^{-t/C})^{2}}.$
Since $\ln(10)>0,C>0$, and the exponential terms are positive, $\psi^{\prime}(t)>0$. Thus, $\psi(t)$ is strictly increasing.
As for antisymmetry, we check if $\psi(-t)=1-\psi(t)$.
$1-\psi(t)$ $=1-\frac{1}{1+10^{-t/C}}=\frac{10^{-t/C}}{1+10^{-t/C}}.$
Multiplying the numerator and denominator by $10^{t/C}$:
$1-\psi(t)=\frac{1}{10^{t/C}+1}=\psi(-t).$
Finally, as $t\to+\infty$, $t/C\to-\infty$, so $10^{-t/C}\to 0$, and $\psi(t)\to 1$. As $t\to$-$\infty$, $t/C\to+\infty$, so $10^{-t/C}\to\infty$, and $\psi(t)\to 0$. ∎
A related concept in Elo systems is the Tournament Performance Rating (TPR), which is a specific type of performance rating function. Given opponents’ pre-tournament ratings, the TPRi of player $i$ is defined as the rating level that would have predicted the observed score $m_{i}$ with respect to $i$’s opponents’ pre-tournament ratings, as opposed to the PRE.
Definition 4 (Tournament Performance Rating).
Assume that player $i$ plays $k_{i}$ games against opponents with ratings $x^{(1)},\dots,x^{(k_{i})}$. Then, TPRi satisfies:
$m_{i}=\sum_{j=1}^{k_{i}}\psi_{\text{Elo}}(\text{TPR}_{i}-x^{(j)}).$ (3)
As Elo noted (Elo, 1978, p. 12), TPR is undefined for zero or perfect scores. [^6] Ismail (2023) introduces Complete Performance Rating (CPR), which is defined for all scores, including when $m_{i}=0$ or $m_{i}=k_{i}$.
Note also that the PRE condition $E(x^{*})=m$ corresponds to finding $x^{*}$ such that for all $i$:
$m_{i}=\sum_{j\neq i}\frac{1}{1+10^{(x^{*}_{j}-x^{*}_{i})/400}}.$
Thus, unlike TPR, PRE does not depend on the pre-tournament ratings and it corresponds to a fixed point of the performance function.
3 Characterization and uniqueness
We characterize the PRE as the solution to a concave maximization problem. This requires assumptions on the tournament structure and outcomes.
Assumption 1 (Irreducibility).
$W$ is irreducible; i.e., $N$ cannot be partitioned into non-empty disjoint sets $S_{1},S_{2}$ such that $w_{ij}=0$ for all $i\in S_{1},j\in S_{2}$.
Irreducibility means that the tournament graph is strongly connected.
Assumption 2 (Non-dominance).
For every non-empty proper subset $S\subset N$, if the total interaction weight across the partition $W(S):=\sum_{i\in S,\,j\notin S}w_{ij}>0$, then the outcomes across the partition are non-extreme:
$\bar{Y}(S^{c}):=\sum_{i\in S,\,j\notin S}y_{ji}>0\quad\text{and}\quad\bar{Y}(S):=\sum_{i\in S,\,j\notin S}y_{ij}>0.$
Non-dominance means that across any split of players who played each other, neither side achieved a 100% score against the other.
We now construct a potential function $P(x)$. Let $A(t)$ be the antiderivative of $\psi(t)$ normalized such that $A(0)=0$:
$A(t)\;=\;\int_{0}^{t}\psi(s)\,ds.$
The antisymmetry of $\psi$ implies $A(t)$ satisfies A(t)A(t)=tA(t)-A(-t)=t.
We define a symmetric pairwise potential function $G:\mathbb{R}^{2}\to\mathbb{R}$:
$G(a,b):=\beta A\left(\frac{a-b}{\beta}\right)+b.$
$G$ is convex since $A(t)$ is convex, as its derivative $\psi(t)$ is increasing. Indeed, it is symmetric: $G(a,b)=G(b,a)$. $\beta A(\frac{b-a}{\beta})+a. G(a,b)G(b,a)=$$G(b,a)=\beta(A(\frac{a-b}{\beta})-A(\frac{b-a}{\beta}))+b-a$. Let $t=(a-b)/\beta$. The expression becomes $\beta(A(t)-A(-t))-\beta t=\beta t-\beta t=0$.
Its derivatives are $\partial_{a}G(a,b)=\beta\psi(\frac{a-b}{\beta})\frac{1}{\beta}=\psi(\Delta_{ab})$, where $\Delta_{ab}:=(a-b)/\beta$. $\partial_{b}G(a,b)=\beta\psi(\frac{a-b}{\beta})\frac{-1}{\beta}+1=1-\psi(\Delta_{ab})=\psi(\Delta_{ba})$.
Finally, it satisfies $G(a+c,b+c)=G(a,b)+c$.
Next, define the convex regularizer $R(x)$ and the concave potential function $P:\mathbb{R}^{n}\to\mathbb{R}$ as:
$R(x):=\sum_{i<j}w_{ij}\,G(x_{i},x_{j}),$
and
$P(x):=m^{\top}x-R(x),$
respectively.
Proposition 2.
$P(x)$ is concave and translation invariant, $P(x+c\mathbf{1})=P(x)$. A vector $x^{*}$ is a PRE if and only if $\nabla P(x^{*})=0$.
Proof.
We first compute the gradient of $R(x)$. For $k\in N$,
$\frac{\partial R}{\partial x_{k}}$ $=\sum_{j>k}w_{kj}\,\partial_{1}G(x_{k},x_{j})\;+\;\sum_{i<k}w_{ik}\,\partial_{2}G(x_{i},x_{k})$
$=\sum_{j>k}w_{kj}\,\psi\!\big(\Delta_{kj}(x)\big)\;+\;\sum_{i<k}w_{ki}\,\psi\!\big(\Delta_{ki}(x)\big)$
$=\sum_{j\neq k}w_{kj}\,\psi\!\big(\Delta_{kj}(x)\big)\;=\;E_{k}(x).$
Thus $\nabla P(x)=m-E(x)$. In particular, the first–order condition $\nabla P(x^{*})=0$ is exactly the PRE condition $E(x^{*})=m$. Since $P$ is concave (see below), any $x^{*}$ with $\nabla P(x^{*})=0$ is a global maximizer of $P$, and any maximizer satisfies $\nabla P(x^{*})=0$.
Summing the expectations over all players and using $\psi\!\big(\Delta_{ij}(x)\big)+\psi\!\big(\Delta_{ji}(x)\big)=1$ yields
$\sum_{i=1}^{n}E_{i}(x)=\sum_{i<j}w_{ij}\Big(\psi\!\big(\Delta_{ij}(x)\big)+\psi\!\big(\Delta_{ji}(x)\big)\Big)=\sum_{i<j}w_{ij}=M=\sum_{i=1}^{n}m_{i}.$
Hence $\mathbf{1}^{\top}\nabla P(x)=\sum_{i}(m_{i}-E_{i}(x))=0$ for all $x$. In particular, at any maximizer of $P$ over $H=\{x:\sum_{i}x_{i}=0\}$, the Lagrange multiplier must be 0.
The Hessian $\nabla^{2}P(x)$ is the negative Jacobian of $E(x)$, and its quadratic form is
$h^{\top}\nabla^{2}P(x)\,h\;=\;-\frac{1}{\beta}\sum_{i<j}w_{ij}\,\psi^{\prime}\!\big(\Delta_{ij}(x)\big)\,(h_{i}-h_{j})^{2}.$ (4)
Since $w_{ij}\geq 0$ and $\psi^{\prime}(t)>0$ for all $t$, the Hessian is negative semidefinite, so $P$ is concave.
For translation invariance, using $G(a+c,b+c)=G(a,b)+c$ we obtain
$R(x+c\mathbf{1})=\sum_{i<j}w_{ij}\big(G(x_{i},x_{j})+c\big)=R(x)+c\,M,$
and therefore
$P(x+c\mathbf{1})=m^{\top}(x+c\mathbf{1})-R(x+c\mathbf{1})=m^{\top}x+c\sum_{i}m_{i}-(R(x)+cM)=P(x),$
since $\sum_{i}m_{i}=M$. ∎
3.1 Existence
Existence requires that the maximum of $P(x)$ is attained. Since $P(x)$ is translation invariant, we analyze its coercivity on the subspace $H=\{x\in\mathbb{R}^{n}:\sum x_{i}=0\}$. $P|_{H}$ is coercive if $P(x)\to$-$\infty$ as $\|x\|\to\infty$, $x\in H$.
Theorem 1 (Existence).
A finite PRE exists if and only if Non-dominance (Assumption 2) holds within each irreducible component of $W$.
Proof.
Because $P$ is translation-invariant, we restrict to $H=\{x:\sum_{i}x_{i}=0\}$ and study the one-dimensional profiles $t\mapsto P(x_{0}+th)$ along rays with $h\in H\setminus\{0\}$. We (i) rewrite the directional derivative $D_{h}P$ in a pairwise form, (ii) define and compute its recession limit $L(h)$, (iii) deduce that $L(h)<0$ for all $h\neq 0$ implies coercivity and hence existence, and (iv) show that failure of Non-dominance produces a direction with $D_{h}P>0$ for all $t\geq 0$, so no maximizer exists.
Step 1: Pairwise form of the directional derivative. Assume $W$ is irreducible (1); otherwise the argument applies component-wise. [^7] If $W$ is reducible, $P$ decomposes as a sum over components and the analysis applies on each component independently, yielding existence if and only if 2 holds on every component. For $h\in H\setminus\{0\}$, the directional derivative is $D_{h}P(x)=h^{\top}\nabla P(x)=h^{\top}(m-E(x))$. Grouping terms by unordered pairs $\{i,j\}$ gives
$D_{h}P(x)$ $=\sum_{i}h_{i}m_{i}-\sum_{i}h_{i}E_{i}(x)$
$=\sum_{i\neq j}h_{i}y_{ij}-\sum_{i\neq j}h_{i}w_{ij}\psi(\Delta_{ij}(x))$
$=\sum_{i<j}(h_{i}y_{ij}+h_{j}y_{ji})-\sum_{i<j}w_{ij}\big(h_{i}\psi(\Delta_{ij})+h_{j}\psi(\Delta_{ji})\big)$
$=\sum_{i<j}(h_{i}-h_{j})\big(y_{ij}-w_{ij}\psi(\Delta_{ij}(x))\big),$ (5)
using $y_{ij}+y_{ji}=w_{ij}$ and $\psi(\Delta_{ij})+\psi(\Delta_{ji})=1$.
Step 2: The recession limit and coercivity. For a fixed base point $x_{0}\in H$, define the ray $x_{t}:=x_{0}+th$ and the recession limit
$L(h):=\lim_{t\to\infty}D_{h}P(x_{t}).$
This limit exists and is independent of $x_{0}$, since $\Delta_{ij}(x_{t})=\Delta_{ij}(x_{0})+\frac{t}{\beta}(h_{i}-h_{j})$. If $L(h)<0$ for every $h\in H\setminus\{0\}$, then $P|_{H}$ is coercive: indeed, by the definition of $L(h)$ there exist $\varepsilon>0$ and $T<\infty$ with $D_{h}P(x_{t})\leq-\varepsilon$ for all $t\geq T$, hence by the fundamental theorem of calculus along the $C^{1}$ curve $t\mapsto x_{t}$,
$P(x_{t})\leq P(x_{T})-\varepsilon\,(t-T)\xrightarrow[t\to\infty$$\infty$.
Step 3: Compute $L(h)\leq 0$. From (3.1) and $\Delta_{ij}(x_{t})=\Delta_{ij}(x_{0})+\tfrac{t}{\beta}(h_{i}-h_{j})$, as $t\to\infty$: if $h_{i}>h_{j}$, then $\psi(\Delta_{ij}(x_{t}))\to 1$ and the pair $\{i,j\}$ contributes $(h_{i}-h_{j})(y_{ij}-w_{ij})=-(h_{i}-h_{j})y_{ji}$; if $h_{i}<h_{j}$, then $\psi(\Delta_{ij}(x_{t}))\to 0$ and the contribution is $(h_{i}-h_{j})y_{ij}=-(h_{j}-h_{i})y_{ij}$. Equivalently, indexing each unordered pair once by the orientation with $h_{i}>h_{j}$, we obtain
[…]
Step 4: Non-dominance $\Rightarrow$ existence. Assume 2 holds. If there were $h\in H\setminus\{0\}$ with $L(h)=0$, then each summand in the equation above must vanish, so $y_{ji}=0$ whenever $h_{i}>h_{j}$. Let $S=\{i:\,h_{i}=\max_{\ell}h_{\ell}\}$; since $h\in H$ and $h\neq 0$, $S$ is a nonempty proper subset. For every $i\in S$ and $j\notin S$, we have $h_{i}>h_{j}$, hence $y_{ji}=0$ and so $\bar{Y}(S^{c})=\sum_{i\in S,\,j\notin S}y_{ji}=0$. By irreducibility (1), $W(S)>0$ (otherwise $S$ would be disconnected), contradicting Non-dominance (2). Thus $L(h)<0$ for all $h\neq 0$, so by Step 2, $P|_{H}$ is coercive and (since $P$ is continuous) attains a maximizer $x^{*}\in H$. At any maximizer under the linear constraint $\sum_{i}x_{i}=0$, the KKT condition gives $\nabla P(x^{*})=\lambda\mathbf{1}$; using $\mathbf{1}^{\top}\nabla P(x)=0$ for all $x$ (mass conservation), we get $\lambda=0$, hence $\nabla P(x^{*})=0$ and therefore $E(x^{*})=m$.
Step 5: Failure of Non-dominance $\Rightarrow$ nonexistence. If 2 fails, there exists a nonempty proper $S\subset N$ with $W(S)>0$ and, say, a sweep by $S$: $y_{ij}=w_{ij}$ for all $i\in S,\,j\notin S$ (equivalently $y_{ji}=0$). Fix $x_{0}\in H$ and set $x_{t}:=x_{0}+th$, where
$h_{i}=\frac{1}{|S|}\ (i\in S),\qquad h_{j}=-\frac{1}{|S^{c}|}\ (j\notin S).$
Then $h\in H$, $h_{i}-h_{j}=0$ within $S$ and within $S^{c}$, and $h_{i}-h_{j}=\frac{1}{|S|}+\frac{1}{|S^{c}|}>0$ across the cut. Using (3.1) and the identities $y_{ji}=w_{ij}-y_{ij}$ and $\psi(\Delta_{ji})=1-\psi(\Delta_{ij})$, one may index each cross pair by $(i\in S,j\notin S)$ and obtain
$D_{h}P(x_{t})=\sum_{i\in S,\;j\notin S}(h_{i}-h_{j})\Big(y_{ij}-w_{ij}\psi\big(\Delta_{ij}(x_{t})\big)\Big).$
Under the sweep assumption $y_{ij}=w_{ij}$, this becomes
$D_{h}P(x_{t})=\Big(\tfrac{1}{|S|}+\tfrac{1}{|S^{c}|}\Big)\sum_{i\in S,\;j\notin S}w_{ij}\,\Big(1-\psi\big(\Delta_{ij}(x_{t})\big)\Big).$
Since $W(S)=\sum_{i\in S,\,j\notin S}w_{ij}>0$ and $\psi:\mathbb{R}\to(0,1)$, each summand is positive for all finite $t$; hence $D_{h}P(x_{t})>0$ for all $t\geq 0$. Therefore $t\mapsto P(x_{t})$ is strictly increasing on $0,\infty)$, so $P$ has no maximizer on $H$ and no finite PRE exists.
Combining Steps 4 and 5 yields the stated equivalence. ∎
3.2 Uniqueness
Uniqueness depends on the strict concavity of $P(x)$, governed by the null space of the Hessian $\nabla^{2}P(x)$.
Theorem 2 (Uniqueness).
Under Irreducibility (Assumption 1), if a PRE exists, it is unique in $H$. The set of all PREs in $\mathbb{R}^{n}$ is the line $\{x^{*}+c\mathbf{1}:c\in\mathbb{R}\}$.
Proof.
From (4), since $\psi^{\prime}(t)>0$ everywhere, $h^{\top}\nabla^{2}P(x)\,h=0$ if and only if $(h_{i}-h_{j})=0$ whenever $w_{ij}>0$. If $W$ is irreducible (the graph is connected), this implies $h_{i}=h_{j}$ for all $i,j$. Thus, the null space of $\nabla^{2}P(x)$ is exactly $\mathrm{span}(\mathbf{1})$.
For any $h\in H,h\neq 0$, we have $h\notin\mathrm{span}(\mathbf{1})$. Therefore, $h^{\top}\nabla^{2}P(x)h<0$. This means $P|_{H}$ (the restriction of $P$ to $H$) is strictly concave. A strictly concave function can have at most one maximizer. If a PRE exists (Theorem 1), the unique maximizer $x^{*}\in H$ is the unique PRE in $H$. The translation invariance of $P(x)$ (Proposition 2) implies the set of all maximizers in $\mathbb{R}^{n}$ is $\{x^{*}+c\mathbf{1}\}$. ∎